OpenStudy (anonymous):

A(3,4) B(5,-2) How do I find the slope?

OpenStudy (anonymous):

mx+b+y good luck :)

OpenStudy (anonymous):

i mean mx+b=y

OpenStudy (anonymous):

rise over run. so the difference in y values over the difference in x. (-2-4)/(5-2) = -6/3 = -2

OpenStudy (anonymous):

another way of stating this is for every 1 unit on the x axis you go over (to the right), you go down 2 units.

OpenStudy (anonymous):

yo bryanj...u posted a problem with an graph earlier this afternoon...about were did the n come from,,,,sis u figure it out

OpenStudy (anonymous):

*did

OpenStudy (anonymous):

ha yeah, i figured it out. i was just hving a brain failure right then haha.

OpenStudy (anonymous):

lmaoooo yo madre me get brain failure...lol....i was trying to figure it out for an hour...lmaooo

OpenStudy (anonymous):

*made

OpenStudy (anonymous):

LOL. yeah, calculus is fun.

OpenStudy (anonymous):

lol...kind of...only when u know what ur doing...lol